# The radius of statistical efficiency

### v1

@@ D.2 Proof of Theorem 4
Taking the derivative of $\psi(t)$ and using the fundamental theorem of calculus we get
[…]
where the second equality follows by taking the Taylor expansion of $\arcsin(s)$. Factorizing a $t^{2}$ from the series yields
$\psi(t)$ $=1+t^{2}\sum_{n=0}^{\infty}{2n\choose n}\frac{|t|^{{n+1}}}{4^{n}(2n+2)}\leq 1+t^{2}\sum_{n=0}^{\infty}{2n\choose n}\frac{1}{4^{n}(2n+2)}=1+\left(\frac{\pi}{2}-1\right)t^{2}=\phi(t),$
where the inequality follows since $|t|\leq 1$ and the second to last equality evaluates (96) at $t=1.$ This completes the proof. ∎
With these two lemmas at hand, we start the proof of Theorem 5.3. We will first verify (29). To this end, notice that we may write $g_{\Sigma}(u,v)=\operatorname{\mathbb{E}}(y_{1}y_{2})^{2}$, where we define the random variables $y_{1}=\langle x,u\rangle$ and $y_{2}=\langle x,v\rangle$. We compute $\operatorname{\mathbb{E}}y_{1}^{2}=\langle\Sigma u,u\rangle$, $\operatorname{\mathbb{E}}y_{2}^{2}=\langle\Sigma v,v\rangle$, and $\operatorname{\mathbb{E}}y_{1}y_{2}=\langle\Sigma u,v\rangle$. Therefore, Lemma E.1 directly implies
$g_{\Sigma}(u,v)=\langle\Sigma u,u\rangle\langle\Sigma v,v\rangle+2\langle\Sigma u,v\rangle^{2}=\|\Sigma^{1/2}u\|_{2}^{2}\cdot\|\Sigma^{1/2}v\|_{2}^{2}+2\langle\Sigma^{1/2}u,\Sigma^{1/2}v\rangle^{2}$
Consequently, applying the Cauchy-Schwarz inequality yields the two sided bound
@@
Next, we verify (28). Notice that the upper bound follows since
$\min_{v\in\mathbb{S}^{d-1}}h_{\Sigma}(u,v)=\min_{v\in\mathbb{S}^{d-1}}\operatorname{\mathbb{E}}\left[\min\{\langle x,u\rangle^{2},\langle x,v\rangle^{2}\}\right]\leq\min_{v\in\mathbb{S}^{d-1}}\operatorname{\mathbb{E}}\left[\langle x,v\rangle^{2}\right]=\lambda_{\min}(\Sigma).$
@@
To prove the lower bound we will show the slightly stronger statement that for any $u,v\in\mathbb{S}^{d-1}$,
$\big{(}1-\tfrac{2}{\pi}\big{)}\lambda_{\min}(\Sigma)\leq h_{\Sigma}(u,v).$
Recall that $\min\{a,b\}=\frac{a+b}{2}-\frac{|a-b|}{2}$ for any $a,b\in\mathbf{R}$. Therefore, we can write
$h_{\Sigma}(u,v)$ $=\operatorname{\mathbb{E}}_{x\sim N(0,\Sigma)}\left[\frac{\langle x,u\rangle^{2}+\langle x,v\rangle^{2}}{2}-\frac{|\langle x,u\rangle^{2}-\langle x,v\rangle^{2}|}{2}\right]$
@@
$=\frac{u^{\top}\Sigma u+v^{\top}\Sigma v}{2}-\frac{\operatorname{\mathbb{E}}_{x\sim N(0,\Sigma)}|\langle x,u\rangle^{2}-\langle x,v\rangle^{2}|}{2}.$
Next, let’s compute
$|\langle x,u\rangle^{2}-\langle x,v\rangle^{2}|=|\underbrace{\langle x,u+v\rangle}_{y_{1}}\underbrace{\langle x,u-v\rangle}_{y_{2}}|.$
Then we get $\sigma_{1}^{2}:=\operatorname{\mathbb{E}}y_{1}^{2}=\|\Sigma^{1/2}(u+v)\|^{2}$ and $\sigma_{2}^{2}:=\operatorname{\mathbb{E}}y_{2}^{2}=\|\Sigma^{1/2}(u-v)\|^{2}$, and
$\operatorname{\mathbb{E}}y_{1}y_{2}$ $=(u+v)^{\top}\Sigma(u-v)=\rho\sigma_{1}\sigma_{2},$
@@
where $\rho:=\frac{\langle\Sigma^{1/2}(u+v),\Sigma^{1/2}(u-v)\rangle}{\|\Sigma^{1/2}(u+v)\|\|\Sigma^{1/2}(u-v)\|}$. Thus, by Lemma E.1 we get
@@
$\operatorname{\mathbb{E}}|y_{1}y_{2}|$ $=\frac{2}{\pi}\left(\underbrace{\sqrt{1-\rho^{2}}+\rho\arcsin(\rho)}_{=:\psi(\rho)}\right)\sigma_{1}\sigma_{2}.$
Therefore, after the relabeling $\hat{u}=\Sigma^{1/2}u$ and $\hat{v}=\Sigma^{1/2}v$ we deduce
$h_{\Sigma}(u,v)=\frac{\|\hat{u}\|^{2}+\|\hat{v}\|^{2}}{2}-\frac{\psi(\rho)}{\pi}\|\hat{u}+\hat{v}\|\|\hat{u}-\hat{v}\|.$
By Lemma E.2 we have that
$h_{\Sigma}(u,v)\geq\frac{\|\hat{u}\|^{2}+\|\hat{v}\|^{2}}{2}-\left(\left(\frac{1}{2}-\frac{1}{\pi}\right)\rho^{2}+\frac{1}{\pi}\right)\|\hat{u}+\hat{v}\|\|\hat{u}-\hat{v}\|.$ (97)
@@
In what follows we upper bound the second term on the right-hand-sight of this inequality. Without loss of generality we assume that $\|\hat{v}\|\leq\|\hat{u}\|.$ By definition $\rho$ is equal to $\cos\alpha$ where $\alpha$ is the angle between $\hat{u}+\hat{v}$ and $\hat{u}-\hat{v}.$ Thus, by the cosine law we have that
$2\rho\|\hat{u}+\hat{v}\|\|\hat{u}-\hat{v}\|=2\cos(\alpha)\|\hat{u}+\hat{v}\|\|\hat{u}-\hat{v}\|=\|\hat{u}+\hat{v}\|^{2}+\|\hat{u}-\hat{v}\|^{2}-4\|\hat{v}\|^{2}=2(\|\hat{u}\|^{2}-\|\hat{v}\|^{2}),$ (98)
@@
where the last equality follows by the parallelogram law. Similarly, using Young’s inequality
$\|\hat{u}+\hat{v}\|\|\hat{u}-\hat{v}\|\leq\frac{\|\hat{u}+\hat{v}\|^{2}+\|\hat{u}-\hat{v}\|^{2}}{2}\leq\|\hat{u}\|^{2}+\|\hat{v}\|^{2}.$ (99)
Then, applying (98) and (99) yields
$\left(\left(\frac{1}{2}-\frac{1}{\pi}\right)\rho^{2}+\frac{1}{\pi}\right)\|\hat{u}+\hat{v}\|\|\hat{u}-\hat{v}\|$ $\leq\left(\frac{1}{2}-\frac{1}{\pi}\right)\rho\left(\|\hat{u}\|^{2}-\|\hat{v}\|^{2}\right)+\frac{1}{\pi}\left(\|\hat{u}\|^{2}+\|\hat{v}\|^{2}\right)$
$\leq\left(\frac{1}{\pi}+\frac{\rho}{2}-\frac{\rho}{\pi}\right)\|\hat{u}\|^{2}+\left(\frac{1}{\pi}-\frac{\rho}{2}+\frac{\rho}{\pi}\right)\|\hat{v}\|^{2}$
$\leq\frac{1}{2}\|\hat{u}\|^{2}+\left(\frac{2}{\pi}-\frac{1}{2}\right)\|\hat{v}\|^{2}$
@@
where the last inequality follows by adding and subtracting $(2/\pi-1/2)$ to the coefficient of $\|\hat{v}\|$ and noting that $\|\hat{v}\|\leq\|\hat{u}\|$. Combining this inequality with (97) yields
$h_{\Sigma}(u,v)\geq\left(1-\frac{2}{\pi}\right)\|\hat{v}\|^{2}=\left(1-\frac{2}{\pi}\right)v^{\top}\Sigma v\geq\left(1-\frac{2}{\pi}\right)\lambda_{\min}(\Sigma),$
which proves the lower bound.
@@
Set $\Sigma_{1}=\operatorname{\mathbb{E}}_{\mu}xx^{\top}$ and $\Sigma_{2}=\operatorname{\mathbb{E}}_{\nu}xx^{\top}$. Define the manifold of rank one $d_{1}\times d_{2}$ matrices:
=$\mathcal{M}=\big{\{}M\in\mathbf{R}^{d_{1}\times d_{2}}:\operatorname{rank}(A)=1\big{\}}.$
A quick computation yields the expression $\nabla f(M)=\operatorname{\mathbb{E}}\langle M-M_{\star},x_{1}x^{\top}_{2}\rangle$. In particular, equality $\nabla f(M_{\star})=0$ holds and therefore the optimal Lagrange multipliers $\lambda^{\star}$ are zero. Hence the Hessian of the Lagrangian $\nabla^{2}\mathcal{L}(M_{\star},\lambda^{\star})$ coincides with $\nabla^{2}f(M_{\star})$. We now successively compute
@@
$=\operatorname{\mathbb{E}}x^{\top}_{1}\Delta(x_{2}x^{\top}_{2})\Delta^{\top}x_{1}$ (101)
$=\operatorname{\mathbb{E}}_{x_{1}}x^{\top}_{1}\Delta\Sigma_{2}\Delta^{\top}x_{1}$ (102)
$=\operatorname{\mathbb{E}}_{x_{1}}\operatorname{tr}(\Delta\Sigma_{2}\Delta^{\top}x_{1}x^{\top}_{1})$ (103)
@@
In particular, if both $\Sigma_{1}$ and $\Sigma_{2}$ are nonsingular, then $\mu\times\nu$ does not lie in $\mathcal{E}$. We now leverage the tangent structure to get an upper bound and a better lower bound on the quadratic form $\nabla^{2}f(M_{\star})[\Delta,\Delta]$. To this end, standard results (see e.g. [10, Section 7.5]) show that the tangent space to $\mathcal{M}$ at $M_{\star}=\beta_{1\star}\beta_{2\star}^{\top}$ is given by
$T_{\mathcal{M}}(M_{\star})=\big{\{}a\beta_{1\star}\beta_{2\star}^{\top}+u\beta_{2\star}^{\top}+\beta_{1\star}v^{\top}:a\in\mathbf{R},\,u\in\beta_{1\star}^{\perp},\,v\in\beta_{2\star}^{\perp}\big{\}}.$
Let $\Delta=a\beta_{1\star}\beta_{2\star}^{\top}+u\beta_{2\star}^{\top}+\beta_{1\star}v^{\top}\in T_{\mathcal{M}}(M_{\star})$. Without loss of generality we might assume $\|\beta_{1\star}\|=\|\beta_{2\star}\|=1,$ since otherwise we can make the change of variables $a^{\prime}/\|\beta_{1\star}\|\|\beta_{2\star}\|\leftarrow a$, $v^{\prime}/\|\beta_{1}\|\leftarrow v$, and $u^{\prime}/\|\beta_{2}\|\leftarrow u$. Thus, $\|\Delta\|_{F}^{2}=a^{2}+\|v\|^{2}+\|u\|^{2}.$ Observe that we may rewrite (105) as
$\nabla^{2}f(M_{\star})[\Delta,\Delta]$ $=\left\|\Sigma^{1/2}_{1}\left((a\beta_{1\star}+u)\beta_{2\star}^{\top}+\beta_{1\star}v^{\top}\right)\Sigma^{1/2}_{2}\right\|_{F}^{2}$ (107)
$=\left\|\left(\Sigma_{2}^{1/2}\otimes\Sigma^{1/2}_{1}\right)\operatorname{vec}\left((a\beta_{1\star}+u)\beta_{2\star}^{\top}\right)+\left(\Sigma^{1/2}_{2}\otimes\Sigma^{1/2}_{1}\right)\operatorname{vec}\left(\beta_{1\star}v^{\top}\right)\right\|_{F}^{2}.$ (108)
To obtain a bound we leverage the following general claim.
@@
Let $A\in S_{+}^{d}$ be a positive definite matrix. Let $x,y\in{\mathbf{R}}^{d}$ be any pair of orthogonal vectors. Then the estimate holds:
$\|Ax+Ay\|^{2}\geq\frac{2}{\kappa(A)^{2}+1}\left(\|Ax\|^{2}+\|Ay\|^{2}\right).$ (109)
@@
The claim will follow from the following inequality
$|\langle Ax,Ay\rangle|\leq\left(1-\frac{2}{\kappa(A)^{2}+1}\right)\|Ax\|\|Ay\|.$ (110)
We will come back to the proof of (110) but first we show how it implies the result (109). Expanding the square yields
@@
$\geq\|Ax\|^{2}-2\left(1-\frac{2}{\kappa(A)^{2}+1}\right)\|Ax\|\|Ay\|+\|Ay\|^{2}$ (113)
$\geq\|Ax\|^{2}-\left(1-\frac{2}{\kappa(A)^{2}+1}\right)\left(\|Ax\|^{2}+\|Ay\|^{2}\right)+\|Ay\|^{2}$ (114)
$=\frac{2}{\kappa(A)^{2}+1}\left(\|Ax\|^{2}+\|Ay\|^{2}\right),$ (115)
where (113) follows by (110) and (114) is an application of Young’s inequality. So we now focus on proving (110). This is equivalent to finding an upper bound for the following optimization problem
$\max_{x,y:~{}x\perp y}~{}\frac{\langle Ax,Ay\rangle}{\|Ax\|\|Ay\|}=\max_{x,y;x\perp y}\frac{\|A(x+y)\|^{2}-\|A(x-y)\|^{2}}{4\|Ax\|\|Ay\|}$ (116)
where the equality follows from the parallelogram law. Using the same law yields that
$\frac{1}{2}\left(\|A(x+y)\|^{2}+\|A(x-y)\|^{2}\right)=\|Ax\|^{2}+\|Ay\|^{2}.$
We use this constraint to define a relaxation of the problem. We introduce two new variables $z$ and $w$, which intuitively play the role of $x+y$ and $x-y$, respectively. Without loss of generality we can set $\|x\|=\|y\|=1$ since we can divide both sides of (110) by $\|x\|\|y\|$ and therefore from orthogonality of $x$ and $y$ we have $\|x-y\|^{2}=1$ and $\|x+y\|^{2}=2$. Define the constraint sets $\mathcal{Z}=\{(z,w)\mid\|z\|^{2}=\|w\|^{2}=2\}$ and $\mathcal{X}_{w,z}=\{(x,y)\mid\|Ax\|^{2}+\|Ay\|^{2}=\left(\|Az\|^{2}+\|Aw\|^{2}\right)/2\}$. We now successively upper bound
$\max_{x,y;x\perp y}\frac{|\langle Ax,Ay\rangle|}{\|Ax\|\|Ay\|}$ $\leq\max_{(z,w)\in\mathcal{Z}}\max_{(x,y)\in\mathcal{X}_{{w,z}}}\frac{\left|\|Az\|^{2}-\|Aw\|^{2}\right|}{4\|Ax\|\|Ay\|}$ (117)
$=\max_{(z,w)\in\mathcal{Z}}\frac{\left|\|Az\|^{2}-\|Aw\|^{2}\right|}{\|Az\|^{2}+\|Aw\|^{2}}$ (118)
$=\frac{\lambda^{2}_{\max}(A)-\lambda^{2}_{\min}(A)}{\lambda^{2}_{\max}(A)+\lambda^{2}_{\min}(A)}$ (119)
where (118) follows since the function $(a,b)\mapsto ab$ constrained to $a^{2}+b^{2}=c$ is minimized at $(a,b)=(\sqrt{c/2},\sqrt{c/2})$ and (119) follows since $(a,b)\mapsto\frac{|a^{2}-b^{2}|}{a^{2}+b^{2}}$ constrained to the interval $\sqrt{2}[\lambda_{\min}(A),\lambda_{\max}(A)]$ attains a maximum at $(a,b)=\sqrt{2}\left(\lambda_{\max}(A),\lambda_{\min}(A)\right).$ Reordering the terms in (119) proves (110). ∎
@@
We instantiate the claim with $A=\Sigma_{2}^{1/2}\otimes\Sigma_{1}^{1/2}$, $x=\operatorname{vec}\left((a\beta_{1\star}+u)\beta_{2\star}^{\top}\right)$ and $y=\operatorname{vec}\left(\beta_{1\star}v^{\top}\right)$. Note that $x$ and $y$ are orthogonal since
$\langle x,y\rangle=\langle(a\beta_{1\star}+u)\beta_{2\star}^{\top},\beta_{1\star}v^{\top}\rangle=\operatorname{tr}\left(\beta_{2\star}(a\beta_{1\star}+u)^{\top}\beta_{1\star}v^{\top}\right)=\operatorname{tr}\left((a\beta_{1\star}+u)^{\top}\beta_{1\star}v^{\top}\beta_{2\star}\right)=0$
@@
where we used the cyclic invariance of the trace and the fact that $\langle v,\beta_{2\star}\rangle=0.$ Further, $\kappa(A)^{2}=\kappa(\Sigma_{1})\kappa(\Sigma_{2})$ and thus, all together we derive

### v2

@@ D.2 Proof of Theorem 4
Taking the derivative of $\psi$ at $t\in[-1,1]$ and using the fundamental theorem of calculus gives
$\psi(t)=1+\int_{0}^{t}\arcsin(s)ds=1+\sum_{n=0}^{\infty}{2n\choose n}\frac{t^{{2n+2}}}{4^{n}(2n+1)(2n+2)},$ (101)
where the last equality follows by integrating the Taylor expansion of $\arcsin(s)$. Therefore
$\psi(t)$ $=1+t^{2}\sum_{n=0}^{\infty}{2n\choose n}\frac{t^{{2n}}}{4^{n}(2n+1)(2n+2)}$
$\leq 1+t^{2}\sum_{n=0}^{\infty}{2n\choose n}\frac{1}{4^{n}(2n+1)(2n+2)}=1+t^{2}\bigg(\frac{\pi}{2}-1\bigg)=\phi(t),$
where the inequality follows since $|t|\leq 1$ and the penultimate equality follows from evaluating (101) at $t=1$ and noting $\psi(1)=\pi/2$. ∎
With these two lemmas in hand, we start the proof of Theorem 5.3. Let us first verify (31). To this end, let $u,v\in\mathbb{S}^{d-1}$ and notice that we may write $g_{\Sigma}(u,v)=\operatorname{\mathbb{E}}(y_{1}y_{2})^{2}$, where we define the centered random variables $y_{1}:=\langle x,u\rangle$ and $y_{2}:=\langle x,v\rangle$ with $x\sim\mathsf{N}(0,\Sigma)$. We compute $\operatorname{\mathbb{E}}y_{1}^{2}=\langle\Sigma u,u\rangle$, $\operatorname{\mathbb{E}}y_{2}^{2}=\langle\Sigma v,v\rangle$, and $\operatorname{\mathbb{E}}y_{1}y_{2}=\langle\Sigma u,v\rangle$. Therefore, Lemma E.1 directly implies
$g_{\Sigma}(u,v)=\langle\Sigma u,u\rangle\langle\Sigma v,v\rangle+2\langle\Sigma u,v\rangle^{2}=\|\Sigma^{1/2}u\|_{2}^{2}\cdot\|\Sigma^{1/2}v\|_{2}^{2}+2\langle\Sigma^{1/2}u,\Sigma^{1/2}v\rangle^{2}.$
Consequently, applying the Cauchy-Schwarz inequality yields the two-sided bound
@@
Next, we verify (30). Notice that the upper bound in (30) follows trivially since for any $u\in\mathbb{S}^{d-1}$,
$\min_{v\in\mathbb{S}^{d-1}}h_{\Sigma}(u,v)=\min_{v\in\mathbb{S}^{d-1}}{\operatorname{\mathbb{E}}}\big[\langle x,u\rangle^{2}\wedge\langle x,v\rangle^{2}\big]\leq\min_{v\in\mathbb{S}^{d-1}}\operatorname{\mathbb{E}}\langle x,v\rangle^{2}=\lambda_{\min}(\Sigma).$
@@
To prove the lower bound in (30), we will show that for any $u,v\in\mathbb{S}^{d-1}$, we have
$\bigg(1-\frac{2}{\pi}\bigg)\langle\Sigma v,v\rangle\leq h_{\Sigma}(u,v).$
To this end, first recall $\min\{a,b\}=\frac{a+b}{2}-\frac{|a-b|}{2}$ for any $a,b\in\mathbf{R}$. Therefore
@@
$h_{\Sigma}(u,v)$ $=\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathsf{N}(0,\Sigma)}\bigg[\frac{\langle x,u\rangle^{2}+\langle x,v\rangle^{2}}{2}-\frac{|\langle x,u\rangle^{2}-\langle x,v\rangle^{2}|}{2}\bigg]$
$=\frac{\langle\Sigma u,u\rangle+\langle\Sigma v,v\rangle}{2}-\frac{\mathop{\operatorname{\mathbb{E}}}_{x\sim\mathsf{N}(0,\Sigma)}|\langle x,u\rangle^{2}-\langle x,v\rangle^{2}|}{2}.$
Now,
$|\langle x,u\rangle^{2}-\langle x,v\rangle^{2}|=|\underbrace{\langle x,u+v\rangle}_{=:y_{1}}\underbrace{\langle x,u-v\rangle}_{=:y_{2}}|.$
Then $\sigma_{1}^{2}:=\operatorname{\mathbb{E}}y_{1}^{2}=\|\Sigma^{1/2}(u+v)\|^{2}_{2}$ and $\sigma_{2}^{2}:=\operatorname{\mathbb{E}}y_{2}^{2}=\|\Sigma^{1/2}(u-v)\|^{2}_{2}$, and
$\operatorname{\mathbb{E}}y_{1}y_{2}$ $=\langle\Sigma(u+v),u-v\rangle=\rho\sigma_{1}\sigma_{2},$
@@
where $\rho:=\frac{\langle\Sigma^{1/2}(u+v),\Sigma^{1/2}(u-v)\rangle}{\|\Sigma^{1/2}(u+v)\|_{2}\|\Sigma^{1/2}(u-v)\|_{2}}$. Thus, Lemma E.1 implies
@@
$\operatorname{\mathbb{E}}|y_{1}y_{2}|$ $=\frac{2}{\pi}\bigg(\underbrace{\sqrt{1-\rho^{2}}+\rho\arcsin(\rho)}_{=:\psi(\rho)}\bigg)\sigma_{1}\sigma_{2}.$
Therefore, after the relabeling $\hat{u}:=\Sigma^{1/2}u$ and $\hat{v}:=\Sigma^{1/2}v$ we deduce
$h_{\Sigma}(u,v)=\frac{\|\hat{u}\|^{2}_{2}+\|\hat{v}\|^{2}_{2}}{2}-\frac{\psi(\rho)}{\pi}\|\hat{u}+\hat{v}\|_{2}\|\hat{u}-\hat{v}\|_{2}.$
By Lemma E.2, we have
$h_{\Sigma}(u,v)\geq\frac{\|\hat{u}\|^{2}_{2}+\|\hat{v}\|^{2}_{2}}{2}-\bigg(\bigg(\frac{1}{2}-\frac{1}{\pi}\bigg)\rho^{2}+\frac{1}{\pi}\bigg)\|\hat{u}+\hat{v}\|_{2}\|\hat{u}-\hat{v}\|_{2}.$ (102)
@@
To complete the proof, we upper-bound the second term on the right-hand side of this inequality. Without loss of generality, we assume $\|\hat{v}\|_{2}\leq\|\hat{u}\|_{2}.$ By definition, $\rho$ is equal to $\cos(\alpha)$ where $\alpha$ is the angle between $\hat{u}+\hat{v}$ and $\hat{u}-\hat{v}$. Thus, by the law of cosines we have
$2\rho\|\hat{u}+\hat{v}\|_{2}\|\hat{u}-\hat{v}\|_{2}=2\cos(\alpha)\|\hat{u}+\hat{v}\|_{2}\|\hat{u}-\hat{v}\|_{2}=\|\hat{u}+\hat{v}\|^{2}_{2}+\|\hat{u}-\hat{v}\|^{2}_{2}-4\|\hat{v}\|^{2}_{2}=2(\|\hat{u}\|^{2}_{2}-\|\hat{v}\|^{2}_{2}),$ (103)
@@
where the last equality follows by the parallelogram law. Similarly, using Young’s inequality:
$\|\hat{u}+\hat{v}\|_{2}\|\hat{u}-\hat{v}\|_{2}\leq\frac{\|\hat{u}+\hat{v}\|^{2}_{2}+\|\hat{u}-\hat{v}\|^{2}_{2}}{2}=\|\hat{u}\|_{2}^{2}+\|\hat{v}\|^{2}_{2}.$ (104)
Then applying (103) and (104) yields
$\bigg(\bigg(\frac{1}{2}-\frac{1}{\pi}\bigg)\rho^{2}+\frac{1}{\pi}\bigg)\|\hat{u}+\hat{v}\|_{2}\|\hat{u}-\hat{v}\|_{2}$ $\leq\bigg(\frac{1}{2}-\frac{1}{\pi}\bigg)\rho\Big(\|\hat{u}\|^{2}_{2}-\|\hat{v}\|^{2}_{2}\Big)+\frac{1}{\pi}\Big(\|\hat{u}\|^{2}_{2}+\|\hat{v}\|^{2}_{2}\Big)$
$=\bigg(\frac{1}{\pi}+\frac{\rho}{2}-\frac{\rho}{\pi}\bigg)\|\hat{u}\|^{2}_{2}+\bigg(\frac{1}{\pi}-\frac{\rho}{2}+\frac{\rho}{\pi}\bigg)\|\hat{v}\|^{2}_{2}$
$\leq\frac{1}{2}\|\hat{u}\|^{2}_{2}+\bigg(\frac{2}{\pi}-\frac{1}{2}\bigg)\|\hat{v}\|^{2}_{2},$
@@
where the last inequality follows by adding and subtracting $(2/\pi-1/2)$ to the coefficient of $\|\hat{v}\|_{2}$ and noting $\|\hat{v}\|_{2}\leq\|\hat{u}\|_{2}$. Combining this inequality with (102) yields
$h_{\Sigma}(u,v)\geq\bigg(1-\frac{2}{\pi}\bigg)\|\hat{v}\|^{2}_{2}=\bigg(1-\frac{2}{\pi}\bigg)\langle\Sigma v,v\rangle\geq\bigg(1-\frac{2}{\pi}\bigg)\lambda_{\min}(\Sigma),$
which proves the lower bound in (30).
@@
Set $\Sigma_{1}=\operatorname{\mathbb{E}}_{\mu}x_{1}x_{1}^{\top}$ and $\Sigma_{2}=\operatorname{\mathbb{E}}_{\nu}x_{2}x_{2}^{\top}$. Let $\mathcal{M}$ denote the manifold of rank-one $d_{1}\times d_{2}$ matrices:
$\mathcal{M}=\big\{M\in\mathbf{R}^{d_{1}\times d_{2}}\mid\operatorname{rank}(A)=1\big\}.$
A quick computation yields the expression $\nabla f(M)=\operatorname{\mathbb{E}}\,\langle M-M_{\star},x_{1}x^{\top}_{2}\rangle x_{1}x_{2}^{\top}$. In particular, equality $\nabla f(M_{\star})=0$ holds and therefore the optimal Lagrange multipliers $\lambda_{\star}$ are zero. Hence the Hessian of the Lagrangian at $(M_{\star},\lambda_{\star})$ coincides coincides with $\nabla^{2}f(M_{\star})$. We now successively compute
@@
$={\operatorname{\mathbb{E}}}\big[x^{\top}_{1}\Delta(x_{2}x^{\top}_{2})\Delta^{\top}x_{1}\big]$ (106)
$={\operatorname{\mathbb{E}}}_{\mu}\big[x^{\top}_{1}\Delta\Sigma_{2}\Delta^{\top}x_{1}\big]$ (107)
$={\operatorname{\mathbb{E}}}_{\mu}\operatorname{tr}(\Delta\Sigma_{2}\Delta^{\top}x_{1}x^{\top}_{1})$ (108)
@@
In particular, if both $\Sigma_{1}$ and $\Sigma_{2}$ are nonsingular, then $\mu\times\nu$ does not lie in $\mathcal{E}$. We now leverage the tangent structure to get a better lower bound on the quadratic form $\nabla^{2}f(M_{\star})[\Delta,\Delta]$. To this end, standard computations (see, e.g., [10, Section 7.5]) show that the tangent space $\mathcal{T}$} of $\mathcal{M}$ at $M_{\star}=\beta_{1\star}\beta_{2\star}^{\top}$ is given by
${\mathcal{T}}$ $=\big\{a\beta_{1\star}\beta_{2\star}^{\top}+u\beta_{2\star}^{\top}+\beta_{1\star}v^{\top}\mid a\in\mathbf{R},\,u\in\beta_{1\star}^{\perp},\,v\in\beta_{2\star}^{\perp}\big\}$ (112)
$=\big\{w\beta_{2\star}^{\top}+\beta_{1\star}v^{\top}\mid w\in\mathbf{R}^{d_{1}},\,v\in\beta_{2\star}^{\perp}\big\}.$ (113)
Henceforth, let $\Delta=w\beta_{2\star}^{\top}+\beta_{1\star}v^{\top}$ be an arbitrary element of $\mathcal{T}$}, where $w\in\mathbf{R}^{d_{1}}$ and $v\in\beta_{2\star}^{\perp}$. Observe that we may rewrite (110) as
$\nabla^{2}f(M_{\star})[\Delta,\Delta]$ $=\left\|\Sigma^{1/2}_{1}\big(w\beta_{2\star}^{\top}+\beta_{1\star}v^{\top}\big)\Sigma^{1/2}_{2}\right\|_{F}^{2}$ (114)
$=\left\|\big(\Sigma_{2}^{1/2}\otimes\Sigma^{1/2}_{1}\big){\operatorname{vec}}\big(w\beta_{2\star}^{\top}\big)+\big(\Sigma^{1/2}_{2}\otimes\Sigma^{1/2}_{1}\big){\operatorname{vec}}\big(\beta_{1\star}v^{\top}\big)\right\|_{2}^{2}.$ (115)
To obtain the desired lower bound in (34), we leverage the following general claim.
@@
Let $A$ be a positive semidefinite linear operator on a Euclidean space $\mathbf{E}$ and let $x,y\in\mathbf{E}$ be any pair of orthogonal vectors. Then the estimate holds:
$\|Ax+Ay\|^{2}\geq\frac{2}{\kappa(A)^{2}+1}\Big(\|Ax\|^{2}+\|Ay\|^{2}\Big).$ (116)
This bound is tight: equality holds if we take $x=(u_{1}+u_{n})/\sqrt{2}$ and $y=(u_{n}-u_{1})/\sqrt{2}$ with $u_{1}$ and $u_{n}$ orthonormal top and least eigenvectors of $A,$ respectively.
@@
The result is trivial if $A$ is singular, so we may assume that $A$ is positive definite. The claim will follow from the inequality
$|\langle Ax,Ay\rangle|\leq\bigg(1-\frac{2}{\kappa(A)^{2}+1}\bigg)\|Ax\|\|Ay\|.$ (117)
Before establishing (117), let us first show how it implies (116). Expanding the square yields
@@
$\geq\|Ax\|^{2}-2\bigg(1-\frac{2}{\kappa(A)^{2}+1}\bigg)\|Ax\|\|Ay\|+\|Ay\|^{2}$ (120)
$\geq\|Ax\|^{2}-\bigg(1-\frac{2}{\kappa(A)^{2}+1}\bigg)\Big(\|Ax\|^{2}+\|Ay\|^{2}\Big)+\|Ay\|^{2}$ (121)
$=\frac{2}{\kappa(A)^{2}+1}\Big(\|Ax\|^{2}+\|Ay\|^{2}\Big),$ (122)
where (120) follows from (117) and (121) is an application of Young’s inequality. Thus, (117) implies (116), and so it now suffices to establish (117), which by the positive definiteness of $A$ is equivalent to
$|\langle x,A^{2}y\rangle|^{2}\leq\Bigg(\frac{\lambda_{\max}(A^{2})-\lambda_{\min}(A^{2})}{\lambda_{\max}(A^{2})+\lambda_{\min}(A^{2})}\Bigg)^{2}\langle x,A^{2}x\rangle\langle y,A^{2}y\rangle.$ (123)
Since $A^{2}$ is positive definite, (123) holds as an instance of Wielandt’s inequality [37, Section 7.4.12].
Finally, a straightforward computation reveals that equality holds in (116) when $x=(u_{1}+u_{n})/\sqrt{2}$ and $y=(u_{n}-u_{1})/\sqrt{2}$ with $u_{1}$ and $u_{n}$ orthonormal top and least eigenvectors of $A,$ respectively. ∎
@@
We now instantiate the claim with $A=\Sigma_{2}^{1/2}\otimes\Sigma_{1}^{1/2}$, $x={\operatorname{vec}}\big(w\beta_{2\star}^{\top}\big)$, and $y={\operatorname{vec}}\big(\beta_{1\star}v^{\top}\big)$. Note that $x$ and $y$ are orthogonal since
$\langle x,y\rangle=\langle w\beta_{2\star}^{\top},\beta_{1\star}v^{\top}\rangle={\operatorname{tr}}\big(\beta_{2\star}w^{\top}\beta_{1\star}v^{\top}\big)={\operatorname{tr}}\big(w^{\top}\beta_{1\star}v^{\top}\beta_{2\star}\big)=\langle w,\beta_{1\star}\rangle\langle v,\beta_{2\star}\rangle=0,$
@@
where we have leveraged the cyclic invariance of the trace and the equality $\langle v,\beta_{2\star}\rangle=0.$ Further, $\kappa(A)^{2}=\kappa(\Sigma_{1})\kappa(\Sigma_{2})$ and $\|\Delta\|_{F}^{2}=\|w\|^{2}_{2}\|\beta_{2\star}\|_{2}^{2}+\|v\|^{2}_{2}\|\beta_{1\star}\|_{2}^{2}$. Thus, altogether we derive